OpenStudy (anonymous):

If xy+6e^y = 6e, find the value of y" at the point where x = 0

OpenStudy (anonymous):

Implicit differentiation: \[xy+6e^y=6e\\ y+x~y'+6e^y~y'=0\\ y'(x+6e^y)=-y\\ y'=-\frac{y}{x+6e^y}\\ \large y''=-\frac{(x+6e^y)~y'-y\left(1+6e^y~y'\right)}{(x+6e^y)^2}\] Now plug in the given number: \(x=0\) \[\begin{align*}x=0~&\Rightarrow~y=1\\ &\Rightarrow~y'=-\frac{1}{0+6e}=-\frac{1}{6e}\\ &\large\Rightarrow~y''=-\frac{(0+6e)~\left(-\frac{1}{6e}\right)-\left(1+6e~\left(-\frac{1}{6e}\right)\right)}{(0+6e)^2}=\cdots \end{align*}\]